OpenStudy (anonymous):

Need help factoring: 2x^2 − 5x + 2 = 0 I know the factors of 2 are: 1 * 2 and -1 * -2 but neither of those equal -5. Can someone help me out? I already know that the answer is (x - 2)(2x - 1) I just don't understand how to get the answer.

OpenStudy (anonymous):

@haileemackk dont u think it should be -4* -1 because here leading coefficient is 2 not 1

OpenStudy (anonymous):

multiply the coefficient of x^2 which is 2 by the constant which is 2 to give you 4. the factors are: -1 * -4 and 1* 4.

OpenStudy (anonymous):

Oh, so I factor by grouping! Okay, now I get it!

Directrix (directrix):

You need two numbers that multiply to 4 and add to -5. What are they? @haileemackk ----- Multiply to a*c where a = 2 and c = 2. AND add to b where b = -5.

OpenStudy (anonymous):

This is what I've got -1 * -4 = 4 and -1 + 4 = -5 Then I would put -1 and -4 into the trinomial, right? 2x^2 − 5x + 2 = 0 would turn into 2x^2 - 1 - 4 + 2 = 0 ?? Split them in half: x^2 - 4 and -5 + 2 But from here I am unsure.

Directrix (directrix):

When everyone else finishes, I will show you what I mean from beginning to end of the factorization.

OpenStudy (shaik0124):

if u want a to solve again i can